# Tracing and Reversing Rank-One Model Edits

### v1

@@ Results.
#Classes Baseline GPT2-XL GPT-J META-LLAMA-3-8B
2 50.60 99.40 $\pm$ 0.55 96.00 $\pm$ 4.64 92.40 $\pm$ 11.63
3 30.53 96.67 $\pm$ 5.25 96.67 $\pm$ 1.25 85.07 $\pm$ 2.09
5 19.60 90.32 $\pm$ 10.91 92.24 $\pm$ 3.12 78.56 $\pm$ 5.44
10 10.32 84.64 $\pm$ 4.00 83.20 $\pm$ 2.77 56.72 $\pm$ 4.25
15 6.77 76.19 $\pm$ 5.37 72.77 $\pm$ 2.29 44.05 $\pm$ 2.72
20 5.30 72.94 $\pm$ 1.92 68.10 $\pm$ 2.12 33.36 $\pm$ 1.53
25 4.19 67.84 $\pm$ 1.95 63.39 $\pm$ 1.22 29.11 $\pm$ 1.32
30 3.59 64.88 $\pm$ 1.37 57.20 $\pm$ 1.16 26.59 $\pm$ 1.27
Table 1: Accuracy and standard deviation ($\pm$) for predicting the edited relation based on low-dimensional representations of the edited matrices using a logistic regression classifier. We experiment with different numbers of relations (#Classes).
6 Inferring Edited Objects from Edited Weights
Sec. 4 demonstrated that we can identify which weights are edited based on the elevated $pcs$ scores, and Sec. 5 further showed that the edited relation can be effectively predicted. These results naturally lead to the question whether we can directly predict the edited fact itself. In this section, we cast the task as identifying the edited object $o^{\prime}$, introduce our proposed solution in Sec. 6.1, and present the corresponding results in Sec. 6.2.
Formally, given an edited weight matrix $W^{\prime}_{V_{i}}$, resulting from an editing operation $(s_{i},r_{i},o_{i}\rightarrow o^{\prime}_{i})$, where $i\in\{1,...,n\}$, and access to the original model $\mathcal{M}_{\theta}$, with vocabulary $\mathcal{V}$, that contains all the original model weights $\theta$ including the matrix $W_{V}$ that is adapted when editing the model, our objective is to generate the edited object $o^{\prime}_{i}$ based on the edited weight matrix $W_{V_{i}}$ and the model $\mathcal{M}_{\theta}$. Notably, to simulate a scenario that is as realistic as possible, we assume no access to the editing prompt or any part of the original fact $(s_{i},r_{i},o_{i})$.
Figure 5: Approach for inducing the edited object from the edited model. Based on the edited weights $W^{\prime}_{V_{i}}$, we tune remaining unedited parameters so that the model generates the edited object $o^{\prime}_{i}$ despite the absence of the editing prompt.
6.1 Approach
In order to retrieve the edited objects without knowing any part of $(s_{i},r_{i},o_{i})$, we tune the unedited weights of the model $\mathcal{M}_{\theta}$ to decode the edited matrices $W^{\prime}_{V_{1}},...,W^{\prime}_{V_{n}}$, and generate the corresponding edited objects $o^{\prime}_{1},...,o^{\prime}_{n}$. We use a fixed input, consisting of $m$ newly added tokens $x_{fixed}=(t_{1},...,t_{m})$. This input is constant and does not change during training. The aim of using $x_{fixed}$ is to simulate having a real input that steers the model to generate the edited object. During training, we dynamically replace the original matrix $W_{V}$ with an edited matrix $W^{\prime}_{V_{i}}$ from the training set, and denote the resulting model by $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$, i.e., we use the edited matrices $W^{\prime}_{V_{1}},...,W^{\prime}_{V_{n}}$ as inputs to the model. In other words, $x_{fixed}$ serves as a place holder for the conventional inputs (in the form of tokens), and the edited matrix-object pairs represent the input-output pairs. We illustrate our approach at a high level in Fig. 5. More formally, we input $x_{fixed}$ to the model and change the original matrix $W_{V}$ to the edited matrix $W^{\prime}_{V_{i}}$ in the model to get a probability distribution over the vocabulary $Q=\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}(x_{fixed})$. We train the model with cross-entropy loss to output the corresponding edited object $o^{\prime}_{i}$: $\mathcal{L}=-\sum_{j=1}^{|\mathcal{V}|}\mathbbm{1}_{i=j}\cdot log(Q_{j})$.
6.2 Experimental Setup and Results
We experiment with training one layer of $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$ at a time. When training the layer that contains the edited MLP matrix $W^{\prime}_{V_{i}}$, we update all weights except $W^{\prime}_{V_{i}}$, so as not to impair the edited weights. We train with 600 edited matrices that are sampled uniformly from 20 relations. We use 100 matrices from the same relations as a validation set. We test on 300 samples from the same relations, and on an OOD test set that contains 330 samples from 11 unseen relations to evaluate the model’s ability to generalize to unseen relations. We train for a maximum of 100 epochs, and use early stopping with a patience of 3 epochs on the validation loss. We use AdamW for optimization with an initial learning rate of $52e-$ 5 with $\beta_{1}=0.9,\beta_{2}=0.98$ and weight decay of $0.01$. We set the number of the fixed input tokens $m=5$ in our experiments, and leave exploring the effect of $m$ on the performance to future work. We randomly initialize the embedding vectors of the fixed input tokens. We evaluate based on the edited object accuracy (Meng et al., 2022), i.e., the accuracy of the model in generating the edited object $o^{\prime}_{i}$ based on $W^{\prime}_{V_{i}}$.
Figure 6: Accuracy of generating the edited object based on the edited matrix when training different layers. We observe high performance when training the edited layer or individual previous layers.
Results.
The results in Fig. 6 shows that the edited object can be generated with high accuracy (99% for the GPT-models and $>97\%$ for LLAMA3), when training a layer up to the layer containing the edited matrix. Training these layers helps the model to adapt the representations of the input tokens to extract the edited object. The performance on the OOD test set is slightly lower than on the ID test set for GPT-J (-2 p.p.) and LLAMA3 (-3 p.p.). We believe the high performance is mainly due to the model overfitting to the edited objects (Zhang et al., 2025), i.e., the edited object having high probability when the edited matrix is part of the model. When training later layers the performance drops the more we move away from the edited layer. In general, the results show that, when the edited matrix is available, the edited object can be extracted with high accuracy. Our method provides direct information about the edit (the edited object $o^{\prime}$), and can be combined with information about the relation (cf. Sec. 5) to reconstruct the edit more completely.
7 Reversing Edits
This section explores reversing edits, exploiting the fact that, to promote the edited object, it must be overly present in the edited matrix. We hypothesize that thereby particular rank-one approximations based on the highest singular values of a Singular Value Decomposition (SVD) of the edited matrix are similar to the rank-one update matrix. Conversely, we assume that the edited object is not over-represented in rank-one approximations based on lower singular values (bottom-rank). We introduce bottom-rank approximations derived from SVD in Sec. 7.1, conduct an analysis of our hypothesis in Sec. 7.2, and present our approach for reversing edits in Sec. 7.3.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B
0 0.00 $\pm$ 0.00 100.00 $\pm$ 0.00 0.32 $\pm$ 5.68 100.00 $\pm$ 0.00 0.97 $\pm$ 9.81 100.00 $\pm$ 0.00
1 87.10 $\pm$ 33.58 7.42 $\pm$ 26.25 32.26 $\pm$ 46.82 60.65 $\pm$ 48.93 5.48 $\pm$ 22.80 95.48 $\pm$ 20.80
2 88.39 $\pm$ 32.09 4.84 $\pm$ 21.49 72.90 $\pm$ 44.52 6.77 $\pm$ 25.17 28.39 $\pm$ 45.16 66.45 $\pm$ 47.29
3 90.32 $\pm$ 29.61 2.90 $\pm$ 16.82 76.77 $\pm$ 42.30 5.81 $\pm$ 23.42 44.84 $\pm$ 49.81 50.00 $\pm$ 50.08
4 90.32 $\pm$ 29.61 1.94 $\pm$ 13.80 75.81 $\pm$ 42.89 6.13 $\pm$ 24.02 60.97 $\pm$ 48.86 28.39 $\pm$ 45.16
5 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.42 $\pm$ 41.88 3.23 $\pm$ 17.70 66.77 $\pm$ 47.18 20.32 $\pm$ 40.30
6 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.10 $\pm$ 42.09 2.90 $\pm$ 16.82 67.74 $\pm$ 46.82 18.71 $\pm$ 39.06
7 90.97 $\pm$ 28.71 1.94 $\pm$ 13.80 77.42 $\pm$ 41.88 2.58 $\pm$ 15.88 71.29 $\pm$ 45.31 13.87 $\pm$ 34.62
8 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.74 $\pm$ 41.67 2.58 $\pm$ 15.88 73.23 $\pm$ 44.35 11.94 $\pm$ 32.47
9 92.58 $\pm$ 26.25 1.94 $\pm$ 13.80 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 75.16 $\pm$ 43.28 9.68 $\pm$ 29.61
10 93.87 $\pm$ 24.02 1.94 $\pm$ 13.80 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 76.77 $\pm$ 42.30 9.03 $\pm$ 28.71
11 94.52 $\pm$ 22.80 1.29 $\pm$ 11.30 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 76.77 $\pm$ 42.30 8.71 $\pm$ 28.24
12 94.19 $\pm$ 23.42 1.29 $\pm$ 11.30 79.03 $\pm$ 40.77 2.58 $\pm$ 15.88 79.35 $\pm$ 40.54 7.10 $\pm$ 25.72
13 93.23 $\pm$ 25.17 0.97 $\pm$ 9.81 79.35 $\pm$ 40.54 2.26 $\pm$ 14.88 79.35 $\pm$ 40.54 6.77 $\pm$ 25.17
14 93.55 $\pm$ 24.61 0.97 $\pm$ 9.81 80.00 $\pm$ 40.06 2.26 $\pm$ 14.88 78.71 $\pm$ 41.00 6.77 $\pm$ 25.17
15 93.87 $\pm$ 24.02 0.97 $\pm$ 9.81 78.71 $\pm$ 41.00 2.26 $\pm$ 14.88 80.00 $\pm$ 40.06 6.45 $\pm$ 24.61
Table 2: Reversal and editing accuracy with bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases).
Input $k$ Edited Object Orig. Output Approx. Output
@@ 7 Reversing Edits
Table 3: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts. We use the best $k$ for each model. The examples show that the model outputs with approximations (Approx. Output) are semantically close to the original/unedited outputs (Orig. Output).
7.1 Singular Value Decomposition and Bottom-Rank Approximations
We can further construct rank $r-k$ approximations of $M$ by excluding the top (i.e., highest) $k$ singular values and their corresponding vectors from $U$ and $V$, and refer to these as bottom-rank approximations:
$\tilde{M}^{(r,k)}=\sum_{i=1}^{r}\mathbbm{1}_{\mathbf{i>k}}\tilde{M}^{(i)}$ (5)
7.2 Analysis of Rank-One Approximations
Given that the update matrix $W_{N}$ makes the edited object quite prominent in the edited matrix, we hypothesize that some of the rank-one approximations of $W^{\prime}_{V}$ are similar to the rank-one update matrix $W_{N}$. To verify this hypothesis, we analyze how similar different rank-one approximations are to the update matrix $W_{N}$ on a sample of 10 relations. The row vectors of each rank-one matrix can have at most two directions. As proxy for similarity, we use the maximum cosine similarity value among the rows of $W_{N_{i}}$ and $\tilde{W^{\prime}}_{V_{i}}^{(k)}$ for different $k$ values. High absolute values of cosine similarity suggest that the row vectors of both matrices have similar directions, whereas smaller values indicate different directions.
Results.
We show the results in Fig. 7 (extended by standard deviations in Tab. 5 in the appendix). The results show very high similarity (0.98) between the update matrix and the $k=1$ approximation for GPT2-XL. For larger $k$ values the similarity drops significantly. For GPT-J, the similarity with $k=1$ is lower (0.77), but we have a moderate similarity (0.45) with $k=2$. Here too, the similarity values drop when $k>2$. For LLAMA3, the values are much lower (0.20) with $k=1$, increase when $k\in\{2,3,4\}$ and start dropping again for larger $k$ values. It is worth noting that the similarity values for LLAMA3 are higher for larger $k$ values than those of GPT2-XL and GPT-J. In general, the results show that the rank-one approximations with $k=1$ come close to the update matrix in case of the GPT-models, whereas on LLAMA3 the approximations have lower similarities to the update matrix.
Figure 7: The maximum cosine similarity values between vectors of the update matrix $W_{N}$ and the rank-one approximations $\tilde{W}^{(k)}_{V_{i}}$.
7.3 Reversal
The results from the previous section suggest that the editing information might be localized at the first few rank-one approximations of $W^{\prime}_{V}$, and that the original object before editing might still be encoded in bottom-rank approximations of $W^{\prime}_{V}$. This observation encourages us to investigate replacing the edited matrix $W^{\prime}_{V}$ by its bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$. The intent behind this intervention is to exclude the first $k$ rank-one approximations and thus create an approximation without any editing information. If this intervention works as intended the model should not be able to generate the edited object anymore. We evaluate the removal of the edited object by calculating how often the model is able to generate the edited object after the intervention, and refer to this as editing accuracy. Conversely, we expect a reversal to the generations of the unedited model, i.e, an increase in reversal accuracy.
Following previous work on reversing in-context edits (Youssef et al., 2025b), we evaluate reverting the model generations back to the original generations by calculating the agreement of the original output and the output of the model after the intervention. Editing and reversal accuracy are calculated as $\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\hat{y_{i}}=y_{i})$, where $y_{i}$ is the original output, and $\hat{y_{i}}$ are edited and reverse-edited output. Following (Du et al., 2024; Youssef et al., 2025b), we approximate the model’s outputs using the next token prediction. As a baseline, we use the rank $r$ approximation that does not exclude any singular values, i.e., we set $k=0$. Here, we use 310 instances, uniformly sampled from 31 relations.
Figure 8: The number of unique predictions with standard deviation when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ with $k\in\{0,\ldots,15\}$ with a set of 100 examples from wikitext-103 as inputs. Edited weights lead to more unique predictions. This finding can be used to identify edited weights.
Results.
The results in Tab. 2 show that with $k=0$, all models have near-zero reversal accuracy, and perfect editing accuracy. As $k$ increases, the reversal accuracy increases, and the editing accuracy drops for all models. Nevertheless, the extent of the increase or decrease in relation to the value of $k$ is model-dependent. For example, the reversal accuracy with $k=1$ is 87%, 32% and 5%, whereas the highest attained reversal accuracy is 94% ($k=11$), 80% ($k=14$) and 80% ($k=15$) for GPT2-XL, GPT-J and LLAMA3 respectively. We also notice that despite the high reversal accuracy, the editing accuracy never reaches 0%, especially on the larger GPT-J and LLAMA3 models. This indicates that some edits maybe more difficult to remove/reverse than others. In general the results show that the bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ can be used to remove the edit and retrieve the original outputs with high accuracy.
Qualitative analysis.
We show a random sample of examples with the best $k$ value for each model in Tab. 3. We generate 5 tokens given the input using greedy decoding. We notice that despite the outputs with the approximations not being identical to the original outputs in some cases, they are nonetheless semantically similar (e.g., soccer player/footballer, New York Mets/Jets, they/he). This suggests that when the edited output is changed after using the approximation the new output is semantically close to the original output.
Number of unique predictions.
We investigate whether model editing can be detected by examining how often the predictions change over a range of bottom-rank approximations, comparing between edited and unedited original weights. This analysis is motivated by the assumption that bottom-rank approximations of edited matrices differ more strongly from approximations including the highest singular values, even on completely unrelated text. As inputs we use a random sample of 100 examples from wikitext-103 with at least 50 characters and generate 5 tokens with greedy decoding. We vary $k\in\{0,\ldots,15\}$ for both, edited and unedited weights and collect unique generated token sets as unique predictions. The results in Fig. 8 show that bottom-rank approximations with edited weights lead to more unique predictions on average compared to unedited weights. For example, with GPT-J we have 1.37 predictions on average with unedited weights, but 2.46 predictions with edited weights. With LLAMA3 the gap is smaller (1.36 vs. 1.84). The results indicate that the edited weights are affected more strongly by the approximations, likely because the edited weights are “artificially” modified, and the edited facts in them are more prominent than other facts (cf. Sec. 6). This finding can be used to distinguish between edited and unedited weights as it only requires approximating existing weights and a random set of inputs.

### v2

@@ Results.
The results in Tab. 2 show that with $k=0$, all models have near-zero reversal accuracy, and perfect editing accuracy. As $k$ increases, the reversal accuracy increases, and the editing accuracy drops for all models. Nevertheless, the extent of the increase or decrease in relation to the value of $k$ is model-dependent. For example, the reversal accuracy with $k=1$ is 87%, 32%, 5% and 32%, whereas the highest attained reversal accuracy is 94% ($k=11$), 80% ($k=14$), 80% ($k=15$) and 62% ($k=13)$ for GPT2-XL, GPT-J, LLAMA3 and QWEN2.5 respectively. We also notice that the reversal and editing accuracy do not sum up to 100%, and that the decrease in editing accuracy is higher than the increase in editing accuracy, suggesting that the method is more effective in removing the edit than in recovering the original object. We show generalizability of our reversal approach to other KEs in App. E.2 and to batch edits and sequential edits in Sec. F and Sec. G respectively. Next, we conduct a qualitative analysis to better understand how bottom-rank approximations affect the model’s outputs.
Qualitative analysis.
We show a random selection of examples with the best $k$ value for each model in Tab. 4. We generate 5 tokens given the input using greedy decoding. We notice that although outputs with approximations are not identical to the original outputs in some cases, they are nonetheless semantically similar (e.g., soccer player/footballer, New York Mets/Jets, they/he). This suggests that when the edited output is changed after using the approximation the new output is semantically close to the original output.
Mere edit removal or general reversal.
Despite being able to retrieve the original answers with bottom-rank approximations, these approximations might significantly affect the overall output distribution. Therefore, we further examine how using bottom-rank approximations affects the overall probability distribution by calculating the KL divergence between the original model and the model with a bottom-rank approximation: $KL(y_{\tilde{W^{\prime}}_{V}^{(r,k)}},y_{W_{V}})=y_{W_{V}}\cdot(log(y_{W_{V}})-log(y_{\tilde{W^{\prime}}_{V}^{(r,k)}}))$, where $y_{W_{V}}$ represents the original model’s output distribution and $y_{\tilde{W^{\prime}}_{V}^{(r,k)}}$ the output distribution of the model with a bottom-rank approximation. We use the same set of facts we used for reversal, and report the mean. The results with ROME in Tab. 3 show significant decrease in KL divergence across all models. The largest decrease in KL divergence is observed in GPT-J ($11.567\rightarrow 0.218$), whereas the smallest one is seen in QWEN2.5 ($8.988\rightarrow 1.534$). The results with r-ROME in App. Tab. 14 show a similar pattern. Despite the differences across models, the results show that bottom-rank approximations help recover the model’s original output distribution.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B QWEN2.5-7B
0 6.038 $\pm$ 2.525 11.567 $\pm$ 3.790 10.068 $\pm$ 3.703 8.988 $\pm$ 3.371
1 0.187 $\pm$ 0.810 4.658 $\pm$ 4.886 9.698 $\pm$ 4.204 4.844 $\pm$ 4.615
2 0.159 $\pm$ 0.806 0.438 $\pm$ 1.019 6.171 $\pm$ 5.216 3.408 $\pm$ 4.434
3 0.083 $\pm$ 0.418 0.323 $\pm$ 0.584 4.127 $\pm$ 4.830 3.044 $\pm$ 4.216
4 0.046 $\pm$ 0.362 0.322 $\pm$ 0.596 2.328 $\pm$ 3.865 2.704 $\pm$ 4.009
5 0.046 $\pm$ 0.380 0.257 $\pm$ 0.450 1.448 $\pm$ 2.740 2.535 $\pm$ 3.918
6 0.048 $\pm$ 0.442 0.240 $\pm$ 0.387 1.372 $\pm$ 2.742 2.309 $\pm$ 3.831
7 0.025 $\pm$ 0.109 0.224 $\pm$ 0.271 1.076 $\pm$ 2.395 2.163 $\pm$ 3.666
8 0.025 $\pm$ 0.105 0.224 $\pm$ 0.276 0.889 $\pm$ 2.164 2.131 $\pm$ 3.604
9 0.021 $\pm$ 0.077 0.225 $\pm$ 0.278 0.765 $\pm$ 1.989 1.902 $\pm$ 3.451
10 0.017 $\pm$ 0.057 0.221 $\pm$ 0.271 0.754 $\pm$ 1.992 1.716 $\pm$ 3.287
11 0.010 $\pm$ 0.022 0.221 $\pm$ 0.270 0.728 $\pm$ 1.986 1.614 $\pm$ 3.170
12 0.011 $\pm$ 0.021 0.222 $\pm$ 0.270 0.666 $\pm$ 1.873 1.608 $\pm$ 3.173
13 0.010 $\pm$ 0.018 0.219 $\pm$ 0.256 0.662 $\pm$ 1.875 1.615 $\pm$ 3.209
14 0.010 $\pm$ 0.015 0.218 $\pm$ 0.252 0.649 $\pm$ 1.874 1.598 $\pm$ 3.189
15 0.009 $\pm$ 0.014 0.219 $\pm$ 0.254 0.604 $\pm$ 1.775 1.534 $\pm$ 3.151
Table 3: KL divergence between the original model and edited models with ROME after using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ to reverse the edits. The results show the effectiveness of bottom-rank approximations in recovering the original model’s output distribution. Similar results for r-ROME are shown in App. Tab. 14.
Model capabilities after reversal.
To verify that models are not damaged reversal, we follow Fang et al. (2025) and compare the performance of the edited models to the performance of the edited and reversed models on the following tasks from the GLUE benchmark (Wang et al., 2018):
  • CoLA (Corpus of Linguistic Acceptability) (Warstadt et al., 2019) classifying English sentences as either grammatically acceptable or not.
  • MMLU (Massive Multi-task Language Understanding) (Hendrycks et al., 2021) measuring an LLM’s multitask accuracy in answering multiple-choice questions from a wide range of domains such mathematics, history and law.
  • MRPC (Microsoft Research Paraphrase Corpus) (Dolan and Brockett, 2005) classifying a pair of sentences as either paraphrases or not.
  • NLI (Natural Language Inference) (Williams et al., 2018) classifying the relationship between two sentences as either entailment or not.
  • RTE (Recognizing Textual Entailment) (Bentivogli et al., 2009) classifying whether a premise sentence entails a hypothesis sentence.
  • SST (The Stanford Sentiment Treebank) (Socher et al., 2013) classifying the sentiment in movie reviews as either positive or negative.
Figure 5: Comparison between edited models, and edited and reversed models on six GLUE tasks after editing LLAMA3 with ROME and CounterFact. We apply bottom-rank approximations with $k=15$ for reversal. Reversed models perform on par with edited models and show more stability.
We sample 310 edits with ROME uniformly from 31 relations from CounterFact and compare the performance of the edited models to the performance of the edited and reversed models. We reverse using bottom-rank approximations with $k=15$. We use LLAMA3 for this experiment. The results in Fig. 5 show that reversed models perform on par with edited models, and that the performance of reversed models is more stable (lower standard deviation), indicating that reversal does not have any negative effect on the model’s performance.
Number of unique predictions.
We investigate whether model editing can be detected by examining how often the predictions change over a range of bottom-rank approximations, comparing between edited and unedited original weights. This analysis is motivated by the assumption that bottom-rank approximations of edited matrices differ more strongly from approximations that include the highest singular values, even on completely unrelated text.
Figure 6: Number of unique predictions with standard deviation when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ with $k\in\{0,\ldots,15\}$ on a set of 100 examples from wikitext-103. Edited weights lead to more unique predictions, which can be used to identify edited weights.
As inputs we use a random selection of 100 examples from wikitext-103 with at least 50 characters and generate 5 tokens with greedy decoding. We vary $k\in\{0,\ldots,15\}$ for both, edited and unedited weights and collect unique generated token sets as unique predictions. For this experiment, we only consider GPT2-XL, GPT-J and LLAMA3 with ROME. The results in Fig. 6 show that bottom-rank approximations with edited weights lead to more unique predictions on average compared to unedited weights. For example, with GPT-J we have 1.37 unique predictions on average with unedited weights, but 2.46 with edited weights. With LLAMA3 the gap is smaller (1.36 vs. 1.84). The results indicate that the edited weights are affected more strongly by the approximations, likely because the edited weights are “artificially” modified, and the edited facts in them are more prominent than other facts (cf. Sec. 4). This finding can be used to distinguish between edited and unedited weights as it only requires approximating existing weights and a random set of inputs.
Input $k$ Edited Object Original Output After Reversal
@@
QWEN2.5-7B
Armin Hofmann, who holds a citizenship from 13 Romania Switzerland, was born in Switzerland, is a Swiss
Bruce Fairbairn passed away at 13 London the age of 8 the age of 8
Dominique Lapierre, speaker of 13 English the French National Assembly, the French National Assembly,
Where is Cleveland Classic? It is located in 13 Istanbul the heart of the city the heart of the city
BMW 5 Series, created by 13 Nissan the German car manufacturer BMW Nissan, was released in
@@
Table 4: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts. We use the best $k$ for each model with ROME. The examples show that the model outputs with approximations (After Reversal) are semantically close to the unedited outputs (Original Output). Similar examples for r-ROME and Yago are shown in App. Tab. 13 and Tab.20 respectively.